Solve by the substitution method.
7m + n= 12
m - 8n= 18
Help me please. This chapter is just aweful for me I understand hardly anything.

Answer by oscargut(2103)

7m + n= 12 (1st eq)
m - 8n= 18 (2nd eq)
so from 1st eq you have n=12-7m
then you substitute on the 2nd
m-8(12-7m)=18
57m-96=18
57m=114
m=2
n=-2